OpenStudy (anonymous):

what is the value of x if 2x-30 and x+15

OpenStudy (anonymous):

Is their any relation between them?? Write your question properly..

OpenStudy (anonymous):

if 2x-30=x+15 then add 30 to both sides of the equation and you get 2x= x+45 move the x to the other side and you get x=45